# Anomalous and linear holographic hard wall models for light unflavored mesons

### v1

@@ IV.1 Linear AHW for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
with $\chi^{2}$/ndf$=72.96/48=1.52.$ This trajectory is presented in the upper left panel of Fig. 3. Both trajectories can be compared with the one obtained from PDG data shown in Table 2 and are in good agreement with it.
Table 4: Masses in MeV for mesons classified under even or odd $\mathcal{CP}$ with isospin $I=0,1$ and spins $S\geq 1$ from the Linear AHW model, determined from Eq. (36), and the percentual deviations $\delta_{\rm Lin}$ of the masses from this model compared with PDG data. For each family the mass of the state with $J=S=1$ is taken as an input. In the last column we show the total dimension $\Delta_{\rm Lin}$ for each state determined by Eq. (35).
$J^{PC}$ $\mathcal{CP}$ I Meson Linear AHW $\delta_{\rm LAHW}$ $\Delta_{\rm Lin}$
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$2^{++}$ even 1 $a_{2}(1320)$ $1296.75\pm 21$ $1.53\%$ $1.85$
$3^{--}$ even 1 $\rho_{3}(1690)$ $1678.61\pm 30$ $0.60\%$ $3.03$
$4^{++}$ even 1 $a_{4}(1970)$ $1993.84\pm 38$ $1.36\%$ $3.87$
$5^{--}$ even 1 $\rho_{5}(2350)$ $2268.03\pm 44$ $2.66\%$ $4.49$
$6^{++}$ even 1 $a_{6}(2450)$ $2513.71\pm 51$ $2.60\%$ $4.96$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 7.83$ $0\%$ 00
$2^{++}$ even 0 $f_{2}(1270)$ $1294.28\pm 16$ $1.47\%$ $1.76$
$3^{--}$ even 0 $\omega_{3}(1670)$ $1669.09\pm 22$ $0.13\%$ $2.88$
$4^{++}$ even 0 $f_{4}(2050)$ $1978.53\pm 28$ $1.96\%$ $3.67$
$5^{--}$ even 0 $\omega_{5}(2250)$ $2247.67\pm 33$ $0.10\%$ $4.24$
$6^{++}$ even 0 $f_{6}(2510)$ $2488.81\pm 38$ $0.97\%$ $4.67$
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 5.0$ $0\%$ 00
$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1667.30\pm 7.89$ $0.20\%$ $0.48$
$3^{+-}$ odd 1 $b_{3}(2030)$ $2000.60\pm 10.75$ $1.55\%$ $0.65$
$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2281.03\pm 13.31$ $1.38\%$ $0.65$
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 10$ $0\%$ 00
$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1646.87\pm 17$ $1.85\%$ $0.72$
$3^{+-}$ odd 0 $h_{3}(2025)$ $2009.96\pm 24$ $0.74\%$ $1.07$
$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2314.19\pm 31$ $0.59\%$ $1.22$
Figure 2: Regge trajectories (blue lines) for mesons, with masses $m$ expressed in GeV, from the Linear AHW model, with total dimension given by Eq. (35). Red dots are the values obtained from Eq. (36) with uncertainties and the blue lines are obtained by linear regression of these points, for each meson family. Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
Figure 3: Regge trajectories for mesons, with masses $m$ expressed in GeV, from the linear AHW model, with total dimension given by Eq. (35). Red dots are the values obtained from Eq. (36) with uncertainties and the blue lines are the Regge trajectories, obtained by linear regression of these 50 points, for each meson family. Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
IV.2 Linear AHW for mesons with even ${\cal CP}$, $I=0$, and $S\geq 1$
For mesons in the linear AHW model with even ${\cal CP}$, $I=1$, and $S\geq 1$, the parameters in Eq. (35) that minimizes $Q$ are $a_{\rm Lin}=6.67\pm 0.13$ and $\epsilon=0$. So, the total dimension reads
$\Delta_{\rm Lin}=(6.67\pm 0.13)\sqrt{S}-2.67\pm 0.13.$ (40)
The masses obtained by this model, using Eq. (36), are presented with respectives uncertainties in Table 4, taking the mass of the state $\omega(782)$ as an input from PDG. Furthermore, we also present in this table the total dimension and the percentual deviations from PDG data for each mesonic state. The Regge trajectory obtained by linear regression of these masses (first six states) is presented in the upper right panel of Fig. 2 and is given by
$J=(0.89\pm 0.01)M^{2}+0.49\pm 0.02,$ (41)
with $\chi^{2}$/ndf$=10.12/4=2.53.$ The upper right panel of Fig. 3 presents the Regge trajectory for the first 50 sates with even ${\cal CP}$, $I=0$, and $S\geq 1$, and it is given by
$J=(0.8766\pm 0.0003)M^{2}+0.52\pm 0.01,$ (42)
with $\chi^{2}$/ndf$=32.16/48=0.67.$ Both equations can be compared with the one obtained from PDG data, shown in Table 2, and are in good agreement with it.
IV.3 Linear AHW for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$
For mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$ ($b_{1}(1235)$ family), the parameters in Eq. (35) that minimizes $Q$ are $a_{\rm Lin}=3.19\pm 0.03$ and $\epsilon=0.05$, so the total dimension becomes
$\Delta_{\rm Lin}=(3.19\pm 0.03)S^{0.55}+0.81\pm 0.03.$ (43)
The masses obtained in this model, using Eq. (36), are presented in Table 4. The Regge trajectory obtained for the first four points is presented in Fig. 2 and is given by
$J=(0.81\pm 0.01)M^{2}-0.24\pm 0.03,$ (44)
with $\chi^{2}$/ndf$=2.86/2=1.43.$
For 50 points the trajectory is in Fig. 3 and is given by
$J=(0.8549\pm 0.0003)M^{2}-0.42\pm 0.01,$ (45)
with $\chi^{2}$/ndf$=163/48=3.39.$ Both equations, considering the uncertainties, are in agreement with the Regge trajectory obtained from PDG data, shown in Table 2.
IV.4 Linear AHW for mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$
The parameters in Eq. (35) that minimizes $Q$ for meson with mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$ ($h_{1}(1170)$ family) are $a_{\rm Lin}=3.79\pm 0.08$ and $\epsilon=0.04$, so the total dimension is
$\Delta_{\rm Lin}=(3.79\pm 0.08)S^{0.54}+0.21\pm 0.08.$ (46)
The masses obtained in this model, using Eq. (36), are presented in Table 4. The Regge trajectory for the first four points is presented in Fig. 2 and is
$J=(0.7509\pm 0.033)M^{2}-0.03\pm 0.01,$ (47)
with $\chi^{2}$/ndf$=0.22/2=0.11.$ The Fig. 3 presents the Regge trajectory for the first 50 points, given by
$J=(0.7667\pm 0.0003)M^{2}-0.06\pm 0.01,$ (48)
with $\chi^{2}$/ndf$=2.20/48=0.046.$ Both trajectories are in good agreement the one from PDG shown in Table 2.
V Conclusions
@@
In this work, we proposed anomalous and linear HW models for light unflavored mesons. We have shown that the masses and Regge trajectories produced by these models are better than that obtained from the original HW and SW models.
@@
Finally, we expect to succesfully extend the analysis presented here to other particles as heavy and heavy-light mesons, as well as baryons. This is presently under investigation.

### v2

@@ IV Linear AHW for mesons
This trajectory is presented in the upper left panel of Fig. 4. The trajectories from Eqs. (31) and (32) can be compared with the one obtained from PDG data shown in Table 1 and are in good agreement with it. These results are synthesized in Table 6.
[…]
For comparison, we also present in the Appendix VI the Tables 9 and 10, with the results of the minimization of $\chi^{2}$ applied to the LAHW model. As one can see from these tables, the results obtained from this method are very close to the minimization of $Q$, Eq. (23), within this model, shown in Tables 5 and 6.
V Discusssion and Conclusions
@@
In this work, we proposed anomalous and linear HW models for light unflavored mesons. We have shown that the masses and Regge trajectories produced by these models are better than those obtained from the original HW and SW models, shown in Table 2 (see, e.g., Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv ; Colangelo:2008us ). In fact, it is remarkable that the mass percentage deviations predicted by the AHW and LAHW models proposed here for mesons are less than $3\%$ when compared with PDG data.
@@
Since we analyzed four meson families classified under $PC$ parity (even or odd) and isospin $I$ values (0 or 1) starting with the states $\rho(770)$ (even $PC$, $I=1$), $\omega(782)$ (even $PC$ and $I=0$), $b_{1}(1235)$ (odd $PC$ and $I=1$), and $h_{1}(1170)$ (odd $PC$ and $I=0$) with the AHW and LAHW models it is possible to discuss some similarities and differences between these families. First of all, the similarities of the $\rho(770)$ and $\omega(782)$ families are very clear from the comparison of the parameters found by best fit to describe them. For instance, looking at Table 4 one sees that the coefficients $a_{\rm Log}$ for these two families are close to each other. The same can be said about the Regge trajectories equations appearing in this table. The same observation holds when analyzing Table 6, with respect to the parameters $a_{\rm LAHW}$ and $\epsilon$, and the corresponding trajectories. Since these two families share the same even $PC$ parity but different isospins, one can conclude that the dependence on the isospin of these parameters is small. The same conclusion is reached with the families of the $b_{1}(1235)$ and $h_{1}(1170)$ mesons which also have the same odd $PC$ parity, but different isospins.
This discussion could also help us understand the role of the $\epsilon$ parameter in Eq. (28). Since we found $\epsilon=0$ for mesons with even $PC$ parity, while $\epsilon\sim 0.05$ in the case of odd $PC$ parity, it seems that this parameter is sensitive to these quantum numbers. This fact is related to the coefficients $\alpha^{\prime}$ and $\alpha_{0}$ of the linear Regge trajectories ($J=\alpha^{\prime}M^{2}+\alpha_{0}$) found for these mesonic families, in the linear anomalous HW model. In the case of even $PC$ parity, $\epsilon=0$, $\alpha^{\prime}\sim 0.9\,{\rm GeV}^{-2}$ and $\alpha_{0}\sim 0.50$, while for odd $PC$ parity one has $\epsilon\sim 0.05$, $\alpha^{\prime}\sim 0.8\,{\rm GeV}^{-2}$ and $\alpha_{0}<0$. From the physical point of view, one can say that the energy necessary to create a meson state with a given spin $S=J$  is lower for an even $PC$ parity one than the case with odd $PC$ parity with the same spin. This is in agreement with the experimental data from PDG, as can be seen in Table 1. Then, it is clear that the nonzero values of $\epsilon$ are related to an extra energy interaction required to form a mesonic state with odd $PC$ parity.
In the case of the AHW Log model, an analogous discussion might be made looking at the values of the $a_{\rm Log}$ parameter, which is $\sim 2.7$ for the even parity case, while $\sim 0.75$ for the odd case. So one could speculate a relation like $a_{\rm Log}\sim 0.75+2\delta_{PC}$, where $\delta_{PC}$ is zero in the odd case (++-+ - or +-+- +) and 1 in the even case (+++++ + or --- -). A similar relation can also be proposed for the LAHW model with $a_{\rm LAHW}\sim 3.5+3.3\,\delta_{PC}=6.8-3.3\,\delta_{PC}$, which could be written more precisely in terms of the parameter $\epsilon$ as $a_{\rm LAHW}\sim 6.8-72\,\epsilon$.
Finally, we expect to successfully extend the analysis presented here to other particles such as heavy and heavy-light mesons, as well as baryons. This is presently under investigation.
@@ Acknowledgments
VI Appendix: $\chi^{2}$ minimization
$I^{G}(J^{PC})$ Meson AHW Log ($\chi^{2}$) $\delta_{\rm Log}$ $\Delta_{\rm Log}$
$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$1^{-}(2^{++})$ $a_{2}(1320)$ $1308.72\pm 22$ $0.62\%$ $1.91$
$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1679.12\pm 31$ $0.57\%$ $3.03$
$1^{-}(4^{++})$ $a_{4}(1970)$ $1986.39\pm 37$ $0.98\%$ $3.83$
$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2259.66\pm 43$ $3.02\%$ $4.44$
$1^{-}(6^{++})$ $a_{6}(2450)$ $2511.59\pm 47$ $2.51\%$ $4.95$
$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
$0^{+}(2^{++})$ $f_{2}(1270)$ $1303.83\pm 22$ $2.22\%$ $1.82$
$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1668.32\pm 30$ $0.08\%$ $2.88$
$0^{+}(4^{++})$ $f_{4}(2050)$ $1972.01\pm 37$ $2.28\%$ $3.63$
$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2242.93\pm 42$ $0.31\%$ $4.22$
$0^{+}(6^{++})$ $f_{6}(2510)$ $2493.25\pm 46$ $1.15\%$ $4.69$
$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1638.56\pm 27$ $1.92\%$ $0.38$
$1^{+}(3^{+-})$ $b_{3}(2030)$ $1988.45\pm 37$ $2.14\%$ $0.60$
$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2312.88\pm 44$ $2.79\%$ $0.76$
$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1629.93\pm 25$ $0.80\%$ $0.66$
$0^{-}(3^{+-})$ $h_{3}(2025)$ $2003.12\pm 33$ $1.08\%$ $1.04$
$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2339.01\pm 39$ $0.47\%$ $1.32$
Table 7: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the AHW Log model, determined from Eq. (21), and the percentage deviations $\delta_{\rm Log}=\delta_{i}\times 100\%$, with $\delta_{i}$ given by Eq. (22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $\chi^{2}$, Eq. (33). For each family the mass of the state with $J=S=1$ is taken as an input from PDG. In the last column we show the anomalous dimension $\Delta_{\rm Log}$ for each state determined by Eq. (20).
We present here the more standard $\chi^{2}$ minimization as a complement to the $Q$ minimization discussed in section III, and also used in section IV to obtain the best fit of PDG data PDG:2022pth . As a second method to obtain the best fit of PDG data from the AHW models, we use the $\chi^{2}$ definition given by
$\chi^{2}=\sum_{i}\frac{|m_{i}-M_{i}|^{2}}{M_{i}},$ (33)
where $m_{i}$ is the mass obtained by our model and $M_{i}$ is the mass from PDG for a given state. So, according to this method, the best fit is such that $\chi^{2}$ given by (33) is minimum.
$I^{G}(J^{PC})$ Meson $a_{\rm Log}$ AHW (Log) Regge Trajectory (GeV) $R^{2}$
$1^{+}(1^{--})$ $\rho(770)$ $2.76\pm 0.12$ $J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03$ 0.9997
$0^{-}(1^{--})$ $\omega(782)$ $2.62\pm 0.11$ $J=(0.89\pm 0.01)M^{2}+0.48\pm 0.03$ 0.9997
$1^{+}(1^{+-})$ $b_{1}(1235)$ $0.55\pm 0.09$ $J=(0.78\pm 0.02)M^{2}-0.14\pm 0.08$ 0.9985
$0^{-}(1^{+-})$ $h_{1}(1170)$ $0.95\pm 0.07$ $J=(0.73\pm 0.01)M^{2}+0.04\pm 0.05$ 0.9993
Table 8: The Regge trajectory in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the AHW model, Eqs. (20) and (21), using minimization with respect to the parameter $\chi^{2}$, Eq. (33). In the last column, we also show the value of $R^{2}$, defined in Eq. (24), for each trajectory.
The results for the AHW model with logarithm anomalous dimensions, Section III, now minimized by the $\chi^{2}$ method are presented in Tables 7 and 8, and can be compared with the results obtained from the minimization of $Q$ shown in Tables 3 and 4. A comparison of these tables shows that the results of the two minimization methods are very close to each other for each family, and actually coincide for the $h_{1}(1170)$ family.
Analogously, the minimization of $\chi^{2}$ can also be applied to the linear AHW model, discussed in Section IV. The results in this case are displayed in Tables 9 and 10. A comparison of these results with the ones from the $Q$ minimization, presented in Tables 5 and 6, shows that both methods produce similar outputs, an even exactly equal for the $\omega(782)$ family.
$I^{G}(J^{PC})$ Meson Linear AHW ($\chi^{2}$) $\delta_{\rm LAHW}$ $\Delta_{\rm LAHW}$
$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$1^{-}(2^{++})$ $a_{2}(1320)$ $1295.28\pm 20$ $1.64\%$ $1.84$
$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1676.08\pm 29$ $0.75\%$ $3.01$
$1^{-}(4^{++})$ $a_{4}(1970)$ $1990.44\pm 37$ $1.19\%$ $3.85$
$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2263.87\pm 44$ $2.84\%$ $4.47$
$1^{-}(6^{++})$ $a_{6}(2450)$ $2508.87\pm 50$ $2.40\%$ $4.93$
$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
$0^{+}(2^{++})$ $f_{2}(1270)$ $1294.28\pm 20$ $1.47\%$ $1.76$
$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1669.09\pm 29$ $0.13\%$ $2.88$
$0^{+}(4^{++})$ $f_{4}(2050)$ $1978.53\pm 37$ $1.96\%$ $3.67$
$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2247.67\pm 44$ $0.10\%$ $4.24$
$0^{+}(6^{++})$ $f_{6}(2510)$ $2488.81\pm 50$ $0.97\%$ $4.67$
$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1665.95\pm 26$ $0.28\%$ $0.48$
$1^{+}(3^{+-})$ $b_{3}(2030)$ $1998.25\pm 36$ $1.66\%$ $0.64$
$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2277.85\pm 45$ $1.24\%$ $0.64$
$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1649.35\pm 26$ $2.00\%$ $0.73$
$0^{-}(3^{+-})$ $h_{3}(2025)$ $2014.27\pm 36$ $0.53\%$ $1.09$
$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2320.03\pm 45$ $0.34\%$ $1.24$
Table 9: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the Linear AHW model, determined from Eq. (29), and the percentage deviations $\delta_{\rm LAHW}=\delta_{i}\times 100\%$, with $\delta_{i}$ from Eq. (22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $\chi^{2}$, Eq. (33). For each family the mass of the state with $J=S=1$ is taken as an input. In the last column we show the total dimension $\Delta_{\rm LAHW}$ for each state determined by Eq. (28).
$I^{G}(J^{PC})$ Meson $a_{\rm LAHW}$ $\epsilon$ Linear AHW Regge Trajectory (GeV) $n$ $R^{2}$
$1^{+}(1^{--})$ $\rho(770)$ $6.85\pm 0.16$ $0.00\pm 0.005$ $J=(0.88\pm 0.01)M^{2}+0.51\pm 0.02$ 6 0.999799
$J=(0.8524\pm 0.0002)M^{2}+0.60\pm 0.01$ 50 0.999997
$0^{-}(1^{--})$ $\omega(782)$ $6.67\pm 0.13$ $0.00\pm 0.005$ $J=(0.89\pm 0.01)M^{2}+0.49\pm 0.02$ 6 0.999845
$J=(0.8766\pm 0.0003)M^{2}+0.52\pm 0.01$ 50 0.999995
$1^{+}(1^{+-})$ $b_{1}(1235)$ $3.18\pm 0.11$ $0.05\pm 0.005$ $J=(0.82\pm 0.01)M^{2}-0.25\pm 0.03$ 4 0.999841
$J=(0.8595\pm 0.0003)M^{2}-0.44\pm 0.01$ 50 0.999993
$0^{-}(1^{+-})$ $h_{1}(1170)$ $3.81\pm 0.12$ $0.04\pm 0.005$ $J=(0.75\pm 0.03)M^{2}-0.02\pm 0.01$ 4 0.999964
$J=(0.7599\pm 0.0003)M^{2}-0.04\pm 0.01$ 50 0.999994
Table 10: The Regge trajectories in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the Linear AHW model, Eq. (29), using the minimization with respect to the parameter $\chi^{2}$, Eq. (33). For each meson family we show two Regge trajectories corresponding to the linear regression of two sets of points with dimension $n$. In the last column, we also show the value of $R^{2}$, defined in Eq. (24), for each trajectory.